OpenStudy (anonymous):

Y=x+5 2x+2y=30 system of equation

OpenStudy (anonymous):

plug Y into the equation: 2x + 2(x+5)=30 2x+2x+10=30 4x=20 x=5 plug x into the other equation: Y=5+5 Y=10 (5,10)